# UDPNet: Unleashing Depth-based Priors for Robust Image Dehazing

### v1

@@ 4.3 Loss Functions
where $i$ denotes the index corresponding to the multiple output branches illustrated in Fig. 5. $\hat{\mathbf{I}}$ represents the restored image, while $\mathbf{Y}$ denotes the ground truth. The $P$ and $S$ correspond to the total elements for normalization. The notation $\cdot$$\cdot$] is a concatenation operator, and $\mathcal{R}$ and $\mathcal{I}$ represent the real and imaginary parts obtained through the fast Fourier transform (FFT). The overall loss can be expressed as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{spatial}}+\lambda\mathcal{L}_{\text{frequency}},$ (11)
the balancing hyper-parameter $\lambda$ is empirically set to $0.1$. The parameters of DepthAnything V2 are frozen during the training stage.
4.4 Remarks
Most existing depth-aware dehazing methods adopt simplistic strategies such as directly concatenating depth and RGB inputs or rely on complex auxiliary schemes like joint training with depth estimators or prompt injection. These methods either fail to effectively leverage the structural depth information or introduce additional training complexity, limiting their practical impact. For example, concatenating depth with RGB features at the input stage often results in biasing the network towards low-frequency cues, which diminishes the restoration of high-frequency details. Similarly, approaches that use auxiliary depth estimators or prompt injection schemes increase optimization complexity without guaranteeing meaningful integration of depth features with the image content.
@@ 4.4 Remarks
In contrast, our framework systematically integrates depth priors through two novel modules: the Depth-Guided Attention Module (DGAM) and the Depth Prior Fusion Module (DPFM). DGAM modulates feature representations adaptively based on depth-driven channel attention, thereby ensuring that structural depth cues are enhanced without compromising high-frequency details. The DPFM, on the other hand, performs multi-scale fusion of depth features with visual representations, enabling the network to maintain spatial consistency and geometric awareness across different layers of feature extraction. This approach addresses the fundamental gap in previous methods, effectively combining depth priors with visual information and preserving both structural integrity and fine details.
Experimental results confirm that our approach outperforms existing methods, particularly in dense and nighttime haze conditions, where traditional RGB-based networks often struggle due to the lack of structural cues. The fusion of depth priors not only improves the visual quality of the restored images but also enhances robustness across varying haze intensities and environments. However, it should be noted that the performance of our method is sensitive to the quality of the depth prior, and extreme degradation variations may reduce the effectiveness of depth-based guidance. Nevertheless, our framework still allows for adaptation to other image restoration tasks, such as deblurring and low-light enhancement, which highlights the broader potential of our design in advancing universal image restoration techniques.
5 Experiments
@@
In this section, we conduct extensive experiments to evaluate the effectiveness and generalization of our proposed UDPNet framework. We first introduce the datasets, metrics, and implementation details. Then we perform quantitative and qualitative assessments of the results produced by the proposed UDPNet and other state-of-the-art methods. In the tables, baseline results are marked in blue for clarity. The best and second-best results are highlighted in boldface and underlined, respectively.
5.1 Datasets and Metrics
Table 1: Summary of datasets used for training and evaluation.
@@ 5.1 Datasets and Metrics
RESIDE Li et al (2018) Synthetic ITS (indoor), OTS (outdoor), SOTS (test, indoor/outdoor) Train (ITS/OTS), Eval (SOTS)
@@
Haze4K Liu et al (2021a) Synthetic 4000 paired images, more realistic haze synthesis Indoor & Outdoor
NH-HAZE Ancuti et al (2020) Real-world 55 paired hazy/clear images, non-homogeneous haze Real-world dehazing
Dense-Haze Ancuti et al (2019) Real-world 55 paired hazy/clear images, homogeneous & dense haze Real-world dehazing
NHR Zhang et al (2020) Synthetic 16,146 train pairs, 1,794 test pairs Nighttime dehazing
GTA5 Yan et al (2020) Synthetic Large-scale, game-engine rendered Nighttime dehazing
@@
SateHaze1k Huang et al (2020) Synthetic 1,200 hazy/clear pairs, satellite imagery Remote sensing image dehazing
RESIDE Li et al (2018) Synthetic OTS & SOTS-outdoor Dehazing (all-in-one task)
@@
Rain100L Yang et al (2019) Synthetic 100 training & 100 testing images Deraining (all-in-one task)
@@
BSD400 Martin et al (2001) Real-world 400 images from Berkeley segmentation Denoising (all-in-one task)
@@
WED Ma et al (2016) Real-world 4,744 high-quality natural images Denoising (all-in-one task)
@@
GoPro Nah et al (2017) Real-world 3,214 blurry/clear pairs from GoPro videos Deblurring (all-in-one task)
LOL-v1 Wei et al (2018) Real-world 500 low/normal-light pairs Low-light enhancement (all-in-one task)
Datasets. To comprehensively validate the capabilities of the proposed UDPNet, we conduct an evaluation across diverse dehazing benchmarks, as presented in Table 1. For daytime scenes, we employ both synthetic and real‐world datasets. The synthetic benchmarks include RESIDE Li et al (2018), which consists of an indoor training set (ITS), an outdoor training set (OTS) and a synthetic objective test set (SOTS), and a more realistica dataset Haze4K Liu et al (2021a). We train separate models on ITS and OTS using a batch size of 6 and an initial learning rate of $4\times 10^{-4}$, and evaluate them on SOTS‑Indoor and SOTS‑Outdoor, respectively. We further train our network on Haze4K under identical settings. For real‑world evaluation, we test on NH‑HAZE Ancuti et al (2020) and Dense‑Haze Ancuti et al (2019), each containing 55 paired hazy and clear images. NH‑HAZE features non‑homogeneous haze, while Dense‑Haze features homogeneous and dense haze. In these experiments, we set the batch size to 2, the initial learning rate to $2\times 10^{-4}$, and use patch crops of size $800\times 1184$ following the previous method Cui et al (2024a) and patch crops of size $600\times 800$ following the previous method Cui et al (2023b). Due to GPU memory constraints, we only incorporate the DGAM in the real-world experiments. For nighttime dehazing, we leverage two nighttime dehazing datasets (i.e., NHR Zhang et al (2020), which provides 16,146 training pairs and 1,794 test pairs with a batch size of 6, and GTA5 Yan et al (2020)). In addition, we evaluate our approach on the remote sensing dataset SateHaze1k Huang et al (2020), given the critical role of image dehazing in remote sensing tasks. In addition to single image dehazing, we evaluate the proposed UDP (Unleashing Depth-based Priors) paradigm in all-in-one settings (i.e., five-task configurations Jiang et al (2025)), thereby demonstrating the robustness of our proposed paradigm. The task settings follow AdaIR Cui et al (2025b) (dehazing on SOTS Li et al (2018), deraining on Rain100L Yang et al (2019), denoising on BSD400 Martin et al (2001) and WED Ma et al (2016), deblurring on GoPro Nah et al (2017), and low-light enhancement on LOL-v1 Wei et al (2018)).
@@
Evaluation Metrics. Across all datasets, we report Peak Signal‑to‑Noise Ratio (PSNR) and Structural Similarity Index (SSIM) Wang et al (2004) to quantitatively assess dehazing quality. Higher values in both metrics indicate superior restoration quality. For consistency and comparability, we calculate metrics based on identical implementation employed in prior studies Cui et al (2024a, 2025b). On the real-world datasets Ancuti et al (2019, 2020), we further evaluate perceptual similarity using LPIPS Zhang et al (2018), where lower scores indicate better perceptual quality.
5.2 Implementation Details
@@
Our framework is implemented using PyTorch Paszke (2019) and trained on two NVIDIA RTX 3090 GPUs. Following the previous methods Guo et al (2022); Cui et al (2023b), each patch is randomly flipped horizontally for data augmentation with a probability of 0.5. The initial learning rate is gradually reduced to 1e-6 with the cosine annealing Loshchilov and Hutter (2016). Adam ($\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) is used for training. Unless specified otherwise, patch size of 256 × 256 is adopted. We adopt the FSNet Cui et al (2023b), ConvIR Cui et al (2024a), and PoolNet Cui et al (2025a) as our baseline to validate the proposed UDPNet. We follow their original training configurations to ensure a fair comparison. The entire training process lasted for a total of 1000 epochs on the Haze4K, ITS and SateHaze1k datasets, 50 epochs on the OTS dataset, 300 epochs on the NHR dataset, 2000 epochs on the GTA5 dataset, and 5000 epochs on the real-world datasets. For universal image restoration, we follow the settings in Cui et al (2025b). The model is trained for 150 epochs with a batch size of 6. We adopt the L1 loss as the sole optimization objective, and training samples are randomly cropped into patches of size $128\times 128$.
@@
Figure 6: Image dehazing comparisons on the SOTS-indoor Li et al (2018) dataset. Please zoom in on screen for a better view.
Figure 7: Image dehazing comparisons on the SOTS-outdoor Li et al (2018) dataset. Please zoom in on screen for a better view.
@@
Table 2: Quantitative comparison on the SOTS-Indoor and SOTS-Outdoor datasets Li et al (2018).
@@
DCP He et al (2010) TPAMI’10 16.61 0.855 19.14 0.861
GridDehazeNet Liu et al (2019) CVPR’19 32.16 0.984 30.86 0.982
MSBDN Dong et al (2020) CVPR’20 33.67 0.985 33.48 0.982
FFA-Net Qin et al (2020) AAAI’20 36.39 0.989 33.57 0.984
AECR-Net Wu et al (2021) CVPR’21 37.17 0.990 - -
MAXIM-2S Tu et al (2022) CVPR’22 38.11 0.991 34.19 0.985
DeHamer Guo et al (2022) CVPR’22 36.63 0.988 35.18 0.986
PMNet Ye et al (2022) ECCV’22 38.41 0.990 34.74 0.985
DehazeFormer-L Song et al (2023) TIP’23 40.05 0.996 - -
FSNet Cui et al (2023b) TPAMI’23 42.45 0.997 40.40 0.997
MB-TaylorFormer-L Qiu et al (2023) ICCV’23 42.64 0.994 38.09 0.991
FocalNet Cui et al (2023a) ICCV’23 40.82 0.996 37.71 0.995
C2PNet Zheng et al (2023) CVPR’23 42.56 0.995 36.68 0.990
DEA-Net-CR Chen et al (2024) TIP’24 41.31 0.995 36.59 0.990
DCMPNet Zhang et al (2024) CVPR’24 42.18 0.997 36.56 0.993
[…]
GridFormer Wang et al (2024c) IJCV’24 42.34 0.994 - -
ConvIR-B Cui et al (2024a) TPAMI’24 42.72 0.997 39.42 0.996
SFMN Shen et al (2025) TIP’25 41.44 0.995 37.72 0.991
PoolNet-B Cui et al (2025a) TIP’25 42.01 0.997 - -
PGH2Net Su et al (2025) AAAI’25 41.70 0.996 37.52 0.989
MB-TaylorFormerV2-L Jin et al (2025) TPAMI’25 42.84 0.995 39.25 0.992
@@
Denotes the unpaired dehazing method.
Table 3: Quantitative results on the Haze4K Liu et al (2021a) dataset.
Method PSNR↑ SSIM↑ DehazeNet Cai et al (2016) 19.12 0.84 GridDehazeNet Liu et al (2019) 23.29 0.93 MSBDN Dong et al (2020) 22.99 0.85 FFA-Net Qin et al (2020) 26.96 0.95 DMT-Net Liu et al (2021a) 28.53 0.96 PMNet Ye et al (2022) 33.49 0.98 MB-TaylorFormer-L Qiu et al (2023) 34.47 0.99 FSNet Cui et al (2023b) 34.12 0.99 GridFormer Wang et al (2024c) 33.27 0.99 ConvIR-B Cui et al (2024a) 34.15 0.99 DEA-Net-CR Chen et al (2024) 34.25 0.99 MB-TaylorFormerV2-B Jin et al (2025) 34.92 0.99 ConvIR + UDP (Ours) 34.82 0.99 FSNet + UDP (Ours) 35.31 0.99
Figure 8: Image dehazing comparisons on the Haze4K Liu et al (2021a) dataset.
Table 4: Quantitative results on the GTA5 Yan et al (2020) dataset.
Method GS Li et al (2015) MRP Zhang et al (2017) Ancuti et al Ancuti et al (2016) Yan et al Yan et al (2020) CycleGAN Zhu et al (2017) Jin et al Jin et al (2023a) ConvIR-B Cui et al (2024a) PoolNet-B Cui et al (2025a) PoolNet + UDP (Ours) ConvIR + UDP (Ours)
@@
Figure 9: Nighttime image dehazing comparisons on the GTA5 Yan et al (2020) dataset.
Table 5: Quantitative results on the NHR Zhang et al (2020) dataset.
Method NDIM Zhang et al (2014) GS Li et al (2015) MRPF Zhang et al (2017) MRP Zhang et al (2017) OSFD Zhang et al (2020) HCD Wang et al (2024b) FocalNet Cui et al (2023a) Jin et al Jin et al (2023a) FSNet Cui et al (2023b) ConvIR-B Cui et al (2024a) PoolNet-B Cui et al (2025a) FSNet + UDP (Ours) ConvIR + UDP (Ours)
@@
Figure 10: Nighttime image dehazing comparisons on the NHR Zhang et al (2020) dataset.
Table 6: Quantitative results on real-world datasets: Dense-Haze Ancuti et al (2019) and NH-HAZE Ancuti et al (2020).
Method Dense-Haze NH-HAZE
@@
DehazeNet Cai et al (2016) 13.84 0.43 - 16.62 0.52 -
MSBDN Dong et al (2020) 15.37 0.49 - 19.23 0.71 -
FFA-Net Qin et al (2020) 14.39 0.45 - 19.87 0.69 -
AECR-Net Wu et al (2021) 15.80 0.47 - 19.88 0.72 -
DeHamer Guo et al (2022) 16.62 0.56 0.6346 20.66 0.68 0.3837
PMNet Ye et al (2022) 16.79 0.51 - 20.42 0.73 -
MB-TaylorFormer-B Qiu et al (2023) 16.66 0.56 0.6125 - - -
C2PNet Zheng et al (2023) 16.88 0.57 - 20.24 0.69 -
FocalNet Cui et al (2023a) 17.07 0.63 0.6087 20.43 0.79 0.3780
SFNet Cui et al (2023c) 17.46 0.58 0.5689 20.46 0.80 -
FSNet Cui et al (2023b) 17.13 0.65 0.5756 20.55 0.81 0.3624
ConvIR-S Cui et al (2024a) 17.45 0.65 0.6000 20.65 0.80 0.3669
PGH2Net Su et al (2025) 17.02 0.61 - - - -
MB-TaylorFormerV2-B Jin et al (2025) 16.95 0.62 - 20.73 0.70 -
@@
Figure 11: Qualitative comparisons on the Dense-Haze Ancuti et al (2019) dataset and NH-HAZE Ancuti et al (2020) dataset among DeHamer Guo et al (2022), FSNet Cui et al (2023b), ConvIR Cui et al (2024a), and ours.
Table 7: Image Dehazing Comparisons on the Remote Sensing Datasets: SateHaze1k-Thin, SateHaze1k-Moderate, and SateHaze1k-Thick Huang et al (2020)
@@
AOD-Net Li et al (2017) 19.54 0.854 20.10 0.885 15.92 0.731
H2RL-Net Chen et al (2021) 20.91 0.880 22.34 0.906 17.41 0.768
FCFT-Net Li and Chen (2020) 23.59 0.913 22.88 0.927 20.03 0.816
Uformer Wang et al (2022) 22.82 0.907 24.47 0.939 20.36 0.815
C2PNet Zheng et al (2023) 19.62 0.880 24.79 0.940 16.83 0.790
Restormer Zamir et al (2022) 23.08 0.912 24.73 0.933 18.58 0.762
Trinity-Net Chi et al (2023) 21.55 0.884 23.35 0.895 20.97 0.823
UMWTransformer Kulkarni et al (2022) 24.29 0.919 26.65 0.946 20.07 0.825
FocalNet Cui et al (2023a) 24.16 0.916 25.99 0.947 21.69 0.847
ConvIR-S Cui et al (2024a) 25.11 0.978 26.79 0.978 22.65 0.950
PoolNet-S Cui et al (2025a) 25.02 0.979 27.02 0.979 22.73 0.955
@@
Denotes methods that are specially designed for remote
sensing image dehazing.
5.3 Comparisons with SOTA Methods

### v2

@@ IV-C Loss Functions
Following Cui et al. [cui2023image, cui2024revitalizing], we employ a multi-scale dual-domain loss to train UDPNet, where the spatial loss is defined as $\mathcal{L}_{\text{spatial}}=\sum_{i=1}^{3}\frac{1}{P_{i}}\|\hat{\mathbf{I}}_{i}-\mathbf{Y}_{i}\|_{1}$ and the frequency-domain loss is $\mathcal{L}_{\text{frequency}}=\sum_{i=1}^{3}\frac{1}{S_{i}}\|[\mathcal{R}(\hat{\mathbf{I}}_{i}),\mathcal{I}(\hat{\mathbf{I}}_{i})]-[\mathcal{R}(\mathbf{Y}_{i}),\mathcal{I}(\mathbf{Y}_{i})]\|_{1}$, where $i$ denotes the index corresponding to the multiple output branches. $\hat{\mathbf{I}}$ represents the restored image, while $\mathbf{Y}$ denotes the ground truth. The $P$ and $S$ correspond to the total elements for normalization. The notation $\cdot$$\cdot$] is a concatenation operator, and $\mathcal{R}$ and $\mathcal{I}$ represent the real and imaginary parts obtained through the fast Fourier transform.
IV-D Remarks
Many existing depth-aware dehazing methods incorporate depth cues either by direct concatenation with inputs or through complex schemes such as joint training with depth estimators or prompt-based conditioning. While effective in some cases, these strategies do not always ensure a stable and practical integration of depth information into image features.
@@ IV-D Remarks
In contrast, our framework systematically integrates depth priors through two plug-and-play modules. DGAM modulates feature representations adaptively based on depth-driven channel attention, thereby ensuring that structural cues are enhanced, while DPFM performs multi-scale fusion to promote spatial consistency across different representation levels. Together, these modules provide a structured way to inject depth information into the restoration pipeline.
Experimental results confirm that our approach yields consistent improvements over existing methods, particularly in various haze conditions where purely RGB-based models may lack geometric cues. Moreover, the fusion is implemented with residual connections and attention-based modulation, allowing the network to down-weight unreliable depth features and fall back to RGB information when depth cues are inconsistent with evidence. This design ensures graceful improvements rather than catastrophic failure when depth estimation is imperfect, as shown in Table IX. Finally, while this work focuses on dehazing, the proposed depth-aware fusion design is potentially extendable to advancing universal image restoration.
V Experiments
@@
In this section, we conduct extensive experiments to evaluate the effectiveness and generalization of our proposed UDPNet. We first introduce the datasets, metrics, and implementation details. Then we perform quantitative and qualitative assessments. In the tables, the baseline results are marked in blue for clarity. The best and second-best results are highlighted in boldface and underlined, respectively.
V-A Datasets and Metrics
TABLE I: Summary of datasets used for training and evaluation.
@@ V-A Datasets and Metrics
Single image dehazing settings
RESIDE [li2018benchmarking] Synthetic ITS (indoor), OTS (outdoor), SOTS Indoor/Outdoor
@@
Haze4K [liu2021synthetic] Synthetic 4000 paired images, more realistic Indoor/Outdoor
NH-HAZE [ancuti2020nh] Real-world 55 paired images, non-homogeneous Real-world
Dense-Haze [ancuti2019dense] Real-world 55 paired images, dense Real-world
NHR [zhang2020nighttime] Synthetic 16,146 train pairs, 1,794 test pairs Nighttime
GTA5 [yan2020nighttime] Synthetic Large-scale, game-engine rendered Nighttime
@@
SateHaze1k [huang2020single] Synthetic 1,200 hazy/clear pairs, satellite imagery Remote sensing
All-in-one image restoration settings
RESIDE [li2018benchmarking] Synthetic OTS & SOTS-outdoor Dehazing
@@
Rain100L [yang2019joint] Synthetic 100 training & 100 testing images Deraining
@@
BSD400 [martin2001database] Real-world 400 images from Berkeley segmentation Denoising
@@
WED [ma2016waterloo] Real-world 4,744 high-quality natural images Denoising
@@
GoPro [nah2017deep] Real-world 3,214 pairs from GoPro videos Deblurring
LOL-v1 [wei2018deep] Real-world 500 low/normal-light pairs Low-light
Datasets. To comprehensively validate the capabilities of the proposed UDPNet, we conduct an evaluation across diverse dehazing benchmarks, as presented in Table I.
@@
Evaluation Metrics. Across all datasets, we report PSNR and SSIM [wang2004image] to quantitatively assess restoration quality. For consistency and comparability, we calculate metrics based on identical implementation employed in prior studies [cui2024revitalizing, cui2025adair]. On the real-world datasets [ancuti2019dense, ancuti2020nh], we further evaluate perceptual similarity using LPIPS [zhang2018unreasonable].
V-B Implementation Details
@@
Our framework is trained on NVIDIA RTX 3090 GPUs. Following the previous methods [cui2023image], each patch is randomly flipped horizontally for data augmentation. The initial learning rate is gradually reduced to 1e-6 with cosine annealing [loshchilov2016sgdr]. Adam ($\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) is used for training. We adopt the FSNet [cui2023image], ConvIR [cui2024revitalizing], and PoolNet [cui2025exploring] as our baselines. We follow their original training configurations to ensure a fair comparison. Due to the high resolution and patch size used in paired real-world datasets, we report a resource-aware configuration, where depth-guided attention module yields the visible gains under limited training budget. We also evaluate the UDP paradigm in all-in-one settings (i.e., five-task configurations [jiang2024survey]), thereby demonstrating the robustness. We adopt the L1 loss as the sole optimization objective.
@@
Figure 6: Image dehazing comparisons on the SOTS-indoor [li2018benchmarking] and SOTS-outdoor datasets [li2018benchmarking]. Please zoom in for a better view.
@@
TABLE II: Quantitative comparison on the SOTS-Indoor and SOTS-Outdoor datasets [li2018benchmarking].
@@
DCP [he2010single] TPAMI’10 16.61 0.855 19.14 0.861
FFA-Net [qin2020ffa] AAAI’20 36.39 0.989 33.57 0.984
AECR-Net [wu2021contrastive] CVPR’21 37.17 0.990 - -
DeHamer [guo2022image] CVPR’22 36.63 0.988 35.18 0.986
DehazeFormer-L [song2023vision] TIP’23 40.05 0.996 - -
FSNet [cui2023image] TPAMI’23 42.45 0.997 40.40 0.997
MB-TaylorFormer-L [qiu2023mb] ICCV’23 42.64 0.994 38.09 0.991
FocalNet [cui2023focal] ICCV’23 40.82 0.996 37.71 0.995
C2PNet [zheng2023curricular] CVPR’23 42.56 0.995 36.68 0.990
DEA-Net-CR [chen2024dea] TIP’24 41.31 0.995 36.59 0.990
DCMPNet [zhang2024depth] CVPR’24 42.18 0.997 36.56 0.993
GridFormer [wang2024gridformer] IJCV’24 42.34 0.994 - -
ConvIR-B [cui2024revitalizing] TPAMI’24 42.72 0.997 39.42 0.996
SFMN [shen2025spatial] TIP’25 41.44 0.995 37.72 0.991
PoolNet-B [cui2025exploring] TIP’25 42.01 0.997 - -
PGH2Net [su2025prior] AAAI’25 41.70 0.996 37.52 0.989
MB-TaylorFormerV2-L [jin2025mb] TPAMI’25 42.84 0.995 39.25 0.992
@@
TABLE III: Quantitative results on the Haze4K dataset [liu2021synthetic].
Method PSNR↑ SSIM↑
DehazeNet [cai2016dehazenet] 19.12 0.84
GridDehazeNet [liu2019griddehazenet] 23.29 0.93
FFA-Net [qin2020ffa] 26.96 0.95
DMT-Net [liu2021synthetic] 28.53 0.96
PMNet [ye2022perceiving] 33.49 0.98
MB-TaylorFormer-L [qiu2023mb] 34.47 0.99
FSNet [cui2023image] 34.12 0.99
GridFormer [wang2024gridformer] 33.27 0.99
ConvIR-B [cui2024revitalizing] 34.15 0.99
DEA-Net-CR [chen2024dea] 34.25 0.99
MB-TaylorFormerV2-B [jin2025mb] 34.92 0.99
ConvIR + UDP (Ours) 34.82 0.99
FSNet + UDP (Ours) 35.31 0.99
TABLE IV: Quantitative results on nighttime image dehazing datasets: GTA5[yan2020nighttime] and NHR[zhang2020nighttime]
The GTA5 dataset [yan2020nighttime]
GS [li2015nighttime] MRP [zhang2017fast] Ancuti et al [ancuti2016night] Yan et al [yan2020nighttime] CycleGAN [zhu2017unpaired] Jin et al [jin2023enhancing] ConvIR-B [cui2024revitalizing] PoolNet-B [cui2025exploring] PoolNet + UDP ConvIR + UDP
@@
The NHR dataset [zhang2020nighttime]
NDIM [zhang2014nighttime] GS [li2015nighttime] MRPF [zhang2017fast] MRP [zhang2017fast] OSFD [zhang2020nighttime] HCD [wang2024restoring] FocalNet [cui2023focal] Jin et al [jin2023enhancing] FSNet [cui2023image] ConvIR-B [cui2024revitalizing] PoolNet-B [cui2025exploring] FSNet + UDP ConvIR + UDP
@@
Figure 7: Image dehazing comparisons on the Haze4K dataset [liu2021synthetic].
Figure 8: Nighttime image dehazing comparisons on the GTA5 dataset [yan2020nighttime].
Figure 9: Nighttime image dehazing comparisons on the NHR dataset [zhang2020nighttime].
TABLE V: Quantitative results on real-world datasets.
Method Dense-Haze [ancuti2019dense] NH-HAZE [ancuti2020nh]
@@
DehazeNet [cai2016dehazenet] 13.84 0.43 - 16.62 0.52 -
MSBDN [dong2020multi] 15.37 0.49 - 19.23 0.71 -
DeHamer [guo2022image] 16.62 0.56 0.6346 20.66 0.68 0.3837
PMNet [ye2022perceiving] 16.79 0.51 - 20.42 0.73 -
MB-TaylorFormer-B [qiu2023mb] 16.66 0.56 0.6125 - - -
C2PNet [zheng2023curricular] 16.88 0.57 - 20.24 0.69 -
FocalNet [cui2023focal] 17.07 0.63 0.6087 20.43 0.79 0.3780
SFNet [cui2023selective] 17.46 0.58 0.5689 20.46 0.80 -
FSNet [cui2023image] 17.13 0.65 0.5756 20.55 0.81 0.3624
ConvIR-S [cui2024revitalizing] 17.45 0.65 0.6000 20.65 0.80 0.3669
PGH2Net [su2025prior] 17.02 0.61 - - - -
MB-TaylorFormerV2-B [jin2025mb] 16.95 0.62 - 20.73 0.70 -
@@
Figure 10: Qualitative comparisons on the Dense-Haze dataset [ancuti2019dense] and NH-HAZE dataset [ancuti2020nh] among DeHamer [guo2022image], FSNet [cui2023image], ConvIR [cui2024revitalizing], and ours.
TABLE VI: Image Dehazing Comparisons on the Remote Sensing Datasets [huang2020single]
@@
AOD-Net [li2017aod] 19.54 0.854 20.10 0.885 15.92 0.731
H2RL-Net [chen2021hybrid] 20.91 0.880 22.34 0.906 17.41 0.768
FCFT-Net [li2020coarse] 23.59 0.913 22.88 0.927 20.03 0.816
C2PNet [zheng2023curricular] 19.62 0.880 24.79 0.940 16.83 0.790
Restormer [zamir2022restormer] 23.08 0.912 24.73 0.933 18.58 0.762
Trinity-Net [chi2023trinity] 21.55 0.884 23.35 0.895 20.97 0.823
UMWTransformer [kulkarni2022unified] 24.29 0.919 26.65 0.946 20.07 0.825
FocalNet [cui2023focal] 24.16 0.916 25.99 0.947 21.69 0.847
ConvIR-S [cui2024revitalizing] 25.11 0.978 26.79 0.978 22.65 0.950
PoolNet-S [cui2025exploring] 25.02 0.979 27.02 0.979 22.73 0.955
@@
Denotes methods that are designed for remote sensing image dehazing.
V-C Comparisons with SOTA Methods